# Panoramic Out-of-Distribution Segmentation

### v1

@@ Abstract
Panoramic imaging enables capturing $360^{\circ}$ images with an ultra-wide Field-of-View (FoV) for dense omnidirectional perception. However, current panoramic semantic segmentation methods fail to identify outliers, and pinhole Out-of-distribution Segmentation (OoS) models perform unsatisfactorily in the panoramic domain due to background clutter and pixel distortions. To address these issues, we introduce a new task, Panoramic Out-of-distribution Segmentation (PanOoS), achieving OoS for panoramas. Furthermore, we propose the first solution, POS, which adapts to the characteristics of panoramic images through text-guided prompt distribution learning. Specifically, POS integrates a disentanglement strategy designed to materialize the cross-domain generalization capability of CLIP. The proposed Prompt-based Restoration Attention (PRA) optimizes semantic decoding by prompt guidance and self-adaptive correction, while Bilevel Prompt Distribution Learning (BPDL) refines the manifold of per-pixel mask embeddings via semantic prototype supervision. Besides, to compensate for the scarcity of PanOoS datasets, we establish two benchmarks: DenseOoS, which features diverse outliers in complex environments, and QuadOoS, captured by a quadruped robot with a panoramic annular lens system. Extensive experiments demonstrate superior performance of POS, with AuPRC improving by $34.25\%$ and FPR95 decreasing by $21.42\%$ on DenseOoS, outperforming state-of-the-art pinhole-OoS methods. Moreover, POS achieves leading closed-set segmentation capabilities. Code and datasets will be available at https://github.com/MengfeiD/PanOoS.
Index Terms:
Scene Segmentation, Panoramic Images, Out-of-Distribution Segmentation, Scene Understanding.
1 Introduction
Panoramic imaging technology, capable of capturing high-quality $360^{\circ}$ images using optical systems, has gained significant attention cross various omnidirectional vision task, such as dense visual prediction [1, 2, 3], holistic scene understanding [4, 5, 6], panoramic scene segmentation [7, 8, 9], and occlusion-aware seamless perception [10]. Especially, for embodied agents, such as autonomous vehicles and mobile robots, an omnidirectional perception facilitates a comprehensive understanding of the surroundings [11, 12, 13]. However, existing panoramic scene segmentation models are trained to recognize a predefined set of semantic classes (e.g., road, building, car, pedestrian[14, 15, 16, 17], fail to identify outliers. Due to the broader Field of View (FoV) of panoramic images and the nature of long-tail semantic distribution in unconstrained surroundings, objects that do not belong to the predefined classes (e.g., animals, furniture) frequently appear in real-world driving environments, which greatly exacerbates the problem. Therefore, there is an urgent need for a technology capable of identifying outliers not present during training to avoid potential dangers.
@@ 1 Introduction
To address the challenge and enhance comprehensive and safe omnidirectional scene understanding, we introduce a new task, termed the Panoramic Out-of-Distribution Segmentation (PanOoS). Compared with pinhole out-of-distribution segmentation (pinhole-OoS), PanOoS offers a broader FoV beyond the pinhole view, for example, ranging from $95^{\circ}$ to $360^{\circ}$ as depicted in Fig. 1, enabling the detection of outliers and even dangers that are not visible from a pinhole perspective, e.g., the tiger. Yet, panoramic images pose unique challenges in visual understanding tasks due to their broader FoV and complex scene content, which introduce pixel distortions [18, 19] and background clutter, complicating semantic interpretation. These characteristics of the panoramic image lead to feature distribution shifts and loss of semantic information, which impair the model’s ability to capture low-level features and fine details of the image. Moreover, these issues interfere with foreground object recognition, introduces semantic ambiguity in overlapping or visually similar regions, and complicates the distinction between in-distribution and out-of-distribution areas.
Figure 2: Comparison of panoramic semantic and out-of-distribution segmentation paradigms. (a) Panoramic semantic segmentation methods adapt models from pinhole to panoramic through domain adaptation, yet remain limited to in-distribution dense predictions; (b) Pinhole-OoS methods combine semantic segmentation with out-of-distribution (OoD) scoring to detect outliers, but are restricted to narrow pinhole FoV; (c) POS achieves unified semantic and OoD segmentation on panoramic images via prompt distribution learning.
@@
Although there are some effective techniques for panoramic semantic segmentation [8, 11, 20], they are designed for closed-set segmentation and do not include outlier detection. The inability to handle outliers means that these techniques may miss critical objects, leading to incorrect or incomplete scene understanding. This issue significantly limits their practical applicability, particularly in dynamic and unpredictable environments where outliers are frequent. High costs associated with data annotation have also resulted in a scarcity of datasets specifically designed for training panoramic segmentation. As illustrated in Fig. 2, most panoramic segmentation methods typically rely on techniques such as unsupervised domain adaptation [9, 11, 10] or knowledge adaptation [12, 21, 22] to transfer models from the pinhole to the panoramic domain. However, these approaches are not suitable for PanOoS, where the model must not only adapt to the characteristics of the panoramic domain, but also learn to distinguish between in-distribution regions and outliers while ensuring closed-set segmentation performance. Meanwhile, state-of-the-art OoS methods [23, 24, 25] for pinhole images typically involve first training a mask-based segmentation model, e.g., in Mask2Former style [26], and then computing anomaly scores using various outlier detection techniques during fine-tuning or inference. Nevertheless, existing research works primarily focus on narrow-FoV pinhole images, making them ill-equipped to address the aforementioned challenges introduced by broader-FoV in panoramic images (see Fig. 3). Unlike panoramic semantic segmentation and pinhole-OoS tasks, PanOoS uniquely faces a conflict between semantic segmentation and out-of-distribution detection within the mask segmentation model architecture.
@@
To tackle the OoS challenge within the panoramic domain, we propose POS, a novel mask-based framework that integrates prompt distribution learning to enhance pixel-level generalization across complex and distorted panoramic visual domains. (1) Compared to purely visual models, vision-language models are typically exposed to a broader array of visual concepts during pre-training [27]. To enhance the pixel-level perception of outliers within panoramic scenes, POS incorporates CLIP [28] into a mask-transformer architecture, leveraging its strong generalization capabilities to better capture scene semantics. Moreover, during outlier exposure fine-tuning, POS explicitly disentangles per-pixel mask decoding from semantic classification, thereby mitigating semantic entanglement and preserving accurate mask representations for subsequent pixel-level decoding. (2) Traditional image-based query decoding methods [26, 23, 24, 25] often struggle to accurately capture semantic information under pixel distortions and background clutter. To address this limitation, we propose a Prompt-based Restoration Attention (PRA), which integrates the text encoder to extract both class-level and distributional prompts semantic information. By adaptively guiding the learning process, the PRA mechanism optimizes semantic extraction, significantly enhancing the accuracy and robustness of semantic understanding, thereby effectively mitigating semantic ambiguity in challenging panoramic scenes. (3) Finally, an innovative Bilevel Prompt Distribution Learning (BPDL) mechanism is proposed. By leveraging the language-image semantic distribution consistency of the CLIP model, BPDL refines the spatial distribution of per-pixel mask feature representations, strengthens the distinguishability of out-of-distribution features, and the adaptability of the pinhole-trained model to severe distortions encountered in panoramic images. With these crucial designs, POS achieves superior alignment with the characteristics of panoramic images, enabling efficient out-of-distribution segmentation while demonstrating impressive closed-set segmentation performance.
@@ 2 Related Work
Recent approaches [62, 29, 30, 63, 64] introduce the Outlier Exposure (OE) strategy [65] that utilizes auxiliary out-of-distribution data to regularize the model’s feature space. These auxiliary outliers typically are images from other datasets (such as ImageNet [66], COCO [67], ADE20K [68], etc.) or involve cutting out abnormal objects from other images and pasting them into inlier scenes. Outliers are used for uncertainty estimation of outlier samples or fine-tuning model parameters. Notably, PEBAL [29] learns energy-based adaptive penalties through abstention learning, whereas DenseHybrid [30] combines likelihood and posterior evaluation to achieve better results. Compared to earlier works, the OE strategy yields better performance, but pixel-wise architecture methods score each pixel individually without considering local semantics, leading to increased noise in anomaly predictions. Currently, mask-transformer-based methods [69, 23, 25, 24, 70, 71] overcome this limitation by segmenting outliers into semantically clustered masks and encouraging the objectness of predictions, which achieve leading segmentation results. EAM [23] ensembles region-based outlier scores, while Mask2Anomaly [25] combines contrastive learning and mask refinement to significantly reduce false positives. RbA [24] utilizes mask queries that act as a one-vs-all classifier [23, 24, 72], helping to eliminate uncertainty in ambiguous in-point regions, e.g., semantic boundaries, by reducing the scores in negative pixels. However, existing methods primarily focus on narrow FoV pinhole images, and when transferred to the panoramic domain, they struggle to adapt to challenges such as background clutter and pixel distortions caused by the broad FoV. To the best of our knowledge, this is the first work to tackle this pressing problem in panoramic out-of-distribution segmentation.
3 Methodology
@@ 3.1 Problem Setting
where $K$ denotes the number of classes for semantic segmentation, $H$ and $W$ correspond to the height and width of the image, respectively.
Figure 4: Overview of POS. The proposed POS consists of four primary components: an image encoder, a text encoder, a pixel decoder, and a transformer decoder with Prompt-based Restoration Attention (PRA). During outlier exposure fine-tuning, Bilevel Prompt Distribution Learning (BPDL) is incorporated to optimize the pixel manifold distribution of the per-pixel mask embeddings. See Fig. 6 for details of BPDL. The image encoder extracts multi-scale panoramic image features in inference, which are then fed into the pixel decoder to generate semantically enriched features alongside the per-pixel embeddings for segmentation. Concurrently, textual prompts are embedded through the text encoder and projected into pixel space. These prompt embeddings, along with the decoded image features, are jointly processed by the transformer decoder, where they interact iteratively with learned query features. The refined queries are subsequently used for mask prediction and category classification. Finally, an out-of-distribution (OoD) scoring mechanism is applied to produce the panoramic OoD segmentation prediction.
3.2 Framework
@@
Text Encoder. For the textual features, the semantic labels with $K$ classes and the distribution labels: inlier and outlier, are integrated into the pre-designed prompt templates, e.g.a photo of a {class.}”, and “a photo of an {outlier} item in the urban or rural scenes”, to obtain the text descriptions about classes and distributions. The text encoder is structured as a $16$-layer transformer, each layer being $768$ units wide and featuring $12$ attention heads. The text descriptions are fed into text encoder to generate the prompt embeddings $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times 768}$, while $T_{0}$ represents the learnable void embedding. Further, a linear projection is used to project the prompt embeddings into the pixel space to match the dimension of the mask feature. See the Appendix A.1 for more details on prompt templates.
@@
Mask and region class prediction. After refining the object queries and mask feature, the refined object queries are processed through a $3$-layer MLP to obtain $Q$ for predicting $N$ regions. Then, the membership score for each pixel belonging to a region is derived by multiplying $Q$ with the per-pixel mask feature $F_{m}(x)$ and applying a $sigmoid$ function:
@@
In parallel, the refined object queries are fed into a linear layer followed by a $softmax$ function to produce posterior class probabilities $P(x){\in}\mathbb{R}^{N\times K}$ for $K$ classes:
@@
$A(x)=RbA(S(x))=-\sum_{k=1}^{K}\phi(P^{T}\!M),$ (4)
with $\phi$ being the $tanh$ activation function. $A(x)$ represents the outlier score map, indicating the outliers in the image.
@@
3.3 Prompt-based Restoration Attention
Figure 5: Prompt-based Restoration Attention continuously employs masked attention and cross attention mechanisms between image features and prompt embeddings while enhancing semantic clarity through query semantic self-adaptive correction.
@@ 3.3 Prompt-based Restoration Attention
$\hat{X}_{i,l}$ $=softmax(\mathcal{M}_{l-1}+Q_{l-1}\leavevmode\nobreak\ K^{T}_{i,l})\leavevmode\nobreak\ V_{i,l},$ (5)
$\hat{X}_{l}$ $=softmax(Q_{i}\leavevmode\nobreak\ K^{T}_{t})\leavevmode\nobreak\ V_{t},$ (6)
@@
Here, LayerNorm (LN) is applied, and $\mathcal{W}$ represents the learnable adaptive weight. We initialize the weight value to zero to ensure the adaptation to diverse multi-scale image features while retaining the training stability. Through the adaptive mechanism, the query dynamically adjusts its representation based on learned semantic information, thereby enhancing semantic generalization.
@@
3.4 Bilevel Prompt Distribution Learning
Figure 6: The demonstration of the process of Bilevel Prompt Distribution Learning. The distribution prompts, text prompts for each class, and pixel embeddings are represented by stars, diamonds, and circles, respectively. The dashed circles indicate equidistant lines from the semantic center, and the dashed lines represent the decision boundary separating in- and out-of-distribution regions.
The pervasive pixel distortions and background clutter in panoramic images often result in a dispersed pixel manifold. This dispersion adversely affects the performance of deep neural networks, which tend to map the unknown samples inside the known class clusters instead of an arbitrary distribution in the whole feature space. Consequently, this misalignment distorts the voting scores of the pixel manifold, particularly in out-of-distribution regions within known classes, ultimately leading to overconfident and unreliable predictions. CLIP, which is trained on large-scale vision-language data pairs, demonstrates robust cross-domain generalization capabilities, enabling it to effectively capture the semantic information of panoramic images. Moreover, its pretrained text encoder adeptly constructs the semantic feature distribution of known classes while maintaining consistency with the pixel space distribution. This accurate semantic feature distribution provides a reliable reference for pixel manifold learning. Leveraging this foundation, we propose a Bilevel Prompt Distribution Learning (DPDL) loss, which utilizes projected class and distribution prompts $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times C_{m}}$, to optimize the spatial distribution of per-pixel embeddings $F_{m}(x)$ for mask segmentation.

### v2

@@ Abstract
Panoramic imaging enables capturing $360^{\circ}$ images with an ultra-wide Field-of-View (FoV) for dense omnidirectional perception, which is critical to applications, such as autonomous driving and augmented reality, etc. However, current panoramic semantic segmentation methods fail to identify outliers, and pinhole Out-of-distribution Segmentation (OoS) models perform unsatisfactorily in the panoramic domain due to background clutter and pixel distortions. To address these issues, we introduce a new task, Panoramic Out-of-distribution Segmentation (PanOoS), with the aim of achieving comprehensive and safe scene understanding. Furthermore, we propose the first solution, POS, which adapts to the characteristics of panoramic images through text-guided prompt distribution learning. Specifically, POS integrates a disentanglement strategy designed to materialize the cross-domain generalization capability of CLIP. The proposed Prompt-based Restoration Attention (PRA) optimizes semantic decoding by prompt guidance and self-adaptive correction, while Bilevel Prompt Distribution Learning (BPDL) refines the manifold of per-pixel mask embeddings via semantic prototype supervision. Besides, to compensate for the scarcity of PanOoS datasets, we establish two benchmarks: DenseOoS, which features diverse outliers in complex environments, and QuadOoS, captured by a quadruped robot with a panoramic annular lens system. Extensive experiments demonstrate superior performance of POS, with AuPRC improving by $34.25\%$ and FPR95 decreasing by $21.42\%$ on DenseOoS, outperforming state-of-the-art pinhole-OoS methods. Moreover, POS achieves leading closed-set segmentation capabilities and advances the development of panoramic understanding. Code and datasets will be available at PanOoS.
I Introduction
Omnidirectional perception technologies play a crucial role in embodied agents, such as autonomous vehicles and mobile robots, facilitating a comprehensive understanding of the surroundings [1, 2]. Especially, panoramic imaging technology, capable of capturing high-quality $360^{\circ}$ images using optical systems, has gained significant attention cross various omnidirectional vision task [3, 4], such as dense visual prediction [5, 6, 7], holistic scene understanding [8, 9, 10], panoramic scene segmentation [11, 12, 13], and occlusion-aware seamless perception [14]. However, existing panoramic scene segmentation models are trained to recognize a predefined set of semantic classes (e.g., road, building, car, pedestrian[15, 16, 17, 18], fail to identify outliers. Due to the broader Field of View (FoV) of panoramic images and the nature of long-tail semantic distribution in unconstrained surroundings, objects that do not belong to the predefined classes (e.g., animals, furniture) frequently appear in real-world driving environments, which greatly exacerbates the problem. Therefore, there is an urgent need for a technology capable of identifying outliers not present during training to avoid potential dangers.
@@ I Introduction
To enhance comprehensive and safe omnidirectional scene understanding and address this challenge, we introduce a new task, termed the Panoramic Out-of-Distribution Segmentation (PanOoS). Compared with pinhole out-of-distribution segmentation (pinhole-OoS), PanOoS offers a broader FoV beyond the pinhole view, for example, ranging from $95^{\circ}$ to $360^{\circ}$ as depicted in Fig. 1, enabling the detection of outliers and even dangers that are not visible from a pinhole perspective, e.g., the tiger. Yet, panoramic images pose unique challenges in visual understanding tasks due to their broader FoV and complex scene content, which introduce pixel distortions [19, 20] and background clutter, complicating semantic interpretation. These characteristics lead to feature distribution shifts and loss of semantic information, which impair the model’s ability to capture low-level features and fine details of the image. Moreover, these issues interfere with foreground object recognition, introduce semantic ambiguity in overlapping or visually similar regions, and complicate the distinction between in-distribution and out-of-distribution areas.
Figure 2: Comparison of panoramic semantic and out-of-distribution segmentation paradigms. (a) Panoramic semantic segmentation methods adapt the model from pinhole to panoramas through domain adaptation, yet remain limited to in-distribution dense predictions; (b) Pinhole-OoS methods combine semantic segmentation with out-of-distribution (OoD) scoring to detect outliers, but are restricted to narrow pinhole FoV; (c) POS achieves segmentation in the distribution space of panoramic images via prompt distribution learning.
@@
Although there are some effective techniques for panoramic semantic segmentation [12, 21, 22], they are only designed for closed-set segmentation. The inability to handle outliers means that these techniques may miss critical objects, leading to incorrect or incomplete scene understanding. This issue significantly limits their practical applicability, particularly in dynamic and unpredictable environments where outliers are frequent. High costs associated with data annotation have also resulted in a scarcity of datasets specifically designed for training panoramic segmentation. As illustrated in Fig. 2, most panoramic segmentation methods typically rely on techniques such as unsupervised domain adaptation [13, 21, 14] or knowledge adaptation [23, 24, 25] to transfer models from the pinhole to the panoramic domain. However, these approaches are not suitable for PanOoS, where the model must not only adapt to the characteristics of the panoramic domain, but also learn to distinguish between in-distribution regions and outliers while ensuring closed-set segmentation performance. Meanwhile, state-of-the-art OoS methods [26, 27, 28] for pinhole images typically involve first training a mask-based segmentation model, e.g., in Mask2Former style [29], and then computing anomaly scores using various outlier detection techniques during fine-tuning or inference. Nevertheless, existing research works primarily focus on narrow-FoV pinhole images, making them ill-equipped to address the aforementioned challenges introduced by broader-FoV in panoramic images (see Fig. 3).
@@
To tackle the OoS challenge within the panoramic domain, we propose POS, a novel mask-based framework that integrates prompt distribution learning to enhance pixel-level generalization across complex and distorted panoramic visual domains. (1) Compared to purely visual models, vision-language models are typically exposed to a broader array of visual concepts during pre-training [30]. To enhance the pixel-level perception of outliers within panoramic scenes, POS incorporates CLIP [31] into a mask-transformer architecture, leveraging its strong generalization capabilities to better capture scene semantics. Moreover, during outlier exposure fine-tuning, POS explicitly disentangles per-pixel mask decoding from semantic classification, thereby mitigating semantic entanglement and preserving accurate mask representations for subsequent pixel-level decoding. (2) Traditional image-based query decoding methods [29, 26, 27, 28] often struggle to accurately capture semantic information under pixel distortions and background clutter. To address this limitation, we propose a Prompt-based Restoration Attention (PRA), which integrates the text encoder to extract both class-level and distributional prompts semantic information. By adaptively guiding the learning process, the PRA mechanism optimizes semantic extraction, significantly enhancing the accuracy and robustness of semantic understanding, thereby effectively mitigating semantic ambiguity in challenging panoramic scenes. (3) Finally, an innovative Bilevel Prompt Distribution Learning (BPDL) mechanism is proposed. By leveraging the language-image semantic distribution consistency of the CLIP model, BPDL refines the spatial distribution of per-pixel mask feature representations and strengthens the distinguishability of out-of-distribution features and the adaptability of the pinhole-trained model to severe distortions encountered in panoramic images. With these crucial designs, POS achieves superior alignment with the characteristics of panoramic images, enabling efficient out-of-distribution segmentation while demonstrating impressive closed-set segmentation performance.
@@ II Related Work
Recent approaches [64, 32, 33, 65, 66] introduce the Outlier Exposure (OE) strategy [67] that utilizes auxiliary out-of-distribution data to regularize the model’s feature space. These auxiliary outliers typically are images from other datasets (such as ImageNet [68], COCO [69], ADE20K [70], etc.) or involve cutting out abnormal objects from other images and pasting them into inlier scenes. Notably, PEBAL [32] learns energy-based adaptive penalties through abstention learning, whereas DenseHybrid [33] combines likelihood and posterior evaluation to achieve better results. Compared to earlier works, the OE strategy yields better performance, but pixel-wise architecture methods score each pixel individually without considering local semantics, leading to increased noise in anomaly predictions. Currently, mask-transformer-based methods [71, 26, 28, 27, 72, 73] overcome this limitation by segmenting outliers into semantically clustered masks and encouraging the objectness of predictions, which achieve leading segmentation results. EAM [26] ensembles region-based outlier scores, while Mask2Anomaly [28] combines contrastive learning and mask refinement to significantly reduce false positives. RbA [27] utilizes mask queries that act as a one-vs-all classifier [26, 27, 74], helping to eliminate uncertainty in ambiguous in-point regions, e.g., semantic boundaries, by reducing the scores in negative pixels. However, existing methods primarily focus on narrow FoV pinhole images, and when transferred to the panoramic domain, they struggle to adapt to challenges such as background clutter and pixel distortions caused by the broad FoV. To the best of our knowledge, this is the first work to tackle this pressing problem in panoramic out-of-distribution segmentation.
III Methodology
@@ III-A Problem Setting
where $K$ denotes the number of classes for closed-set segmentation, $H$ and $W$ correspond to the height and width of the image, respectively.
Figure 4: Overview of POS. The proposed POS consists of four primary components: an image encoder, a text encoder, a pixel decoder, and a transformer decoder with Prompt-based Restoration Attention (PRA). During outlier exposure fine-tuning, Bilevel Prompt Distribution Learning (BPDL) is incorporated to optimize the pixel manifold distribution of the per-pixel mask embeddings. See Fig. 6 for details of BPDL. The image encoder extracts multi-scale panoramic image features in inference, which are then fed into the pixel decoder to generate semantically enriched features alongside the per-pixel embeddings for segmentation. Concurrently, textual prompts are embedded through the text encoder and projected into pixel space. These prompt embeddings, along with the decoded image features, are jointly processed by the transformer decoder, where they interact iteratively with learned query features. The refined queries are subsequently used for mask prediction and category classification. Finally, an OoD scoring mechanism is applied to produce the panoramic OoD segmentation prediction.
III-B Framework
@@
Text Encoder. For the textual features, the semantic labels with $K$ classes and the distribution labels: inlier and outlier, are integrated into the pre-designed prompt templates, e.g.a photo of a {car}”, and “a photo of an {outlier} item in the urban or rural scenes”, to obtain the text descriptions about classes and distributions. The text encoder is structured as a $16$-layer transformer, each layer being $768$ units wide and featuring $12$ attention heads. The text descriptions are fed into text encoder to generate the prompt embeddings $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times 768}$, while $T_{0}$ represents the learnable void embedding. Further, a linear projection is used to project the prompt embeddings into the pixel space to match the dimension of the mask feature. See the Appendix A-A for more details on prompt templates.
@@
Mask and region class prediction. After refining the object queries and mask feature, the refined object queries are processed through a $3$-layer MLP to obtain $Q$ for predicting $N$ regions. Then, the membership score for each pixel belonging to a region is derived by multiplying $Q$ with the per-pixel mask feature $F_{m}(x)$ and applying a sigmoidsigmoid function:
@@
In parallel, the refined object queries are fed into a linear layer followed by a softmaxsoftmax function to produce posterior class probabilities $P(x){\in}\mathbb{R}^{N\times K}$ for $K$ classes:
@@
$A(x)=RbA(S(x))=-\sum_{k=1}^{K}\phi(P(x)^{T}\!M(x)),$ (4)
with $\phi$ being the tanhtanh activation function. $A(x)$ represents the outlier score map, indicating the outliers in the image.
@@
III-C Prompt-based Restoration Attention
@@
$\hat{X}_{i,l}$ $=softmax(\mathcal{M}_{l-1}+Q_{l-1}~K^{T}_{i,l})~V_{i,l},$ (5)
$\hat{X}_{l}$ $=softmax(Q_{i}~K^{T}_{t})~V_{t},$ (6)
@@
Figure 5: Prompt-based Restoration Attention continuously employs masked attention and cross attention between image features and prompt embeddings while enhancing semantic clarity through query semantic self-adaptive correction.
@@
Here, LayerNorm (LN) is applied, and $\mathcal{W}$ represents the learnable adaptive weight. We initialize the weight value to zero to ensure the adaptation to diverse multi-scale image features while retaining the training stability. Through the adaptive mechanism, the query dynamically adjusts its representation based on learned semantic information.
@@
III-D Bilevel Prompt Distribution Learning
Figure 6: Bilevel Prompt Distribution Learning. The distribution prompts, text prompts for each class, and pixel embeddings are represented by stars, diamonds, and circles, respectively. The dashed circles indicate equidistant lines from the semantic center, and the dashed lines represent the decision boundary separating in- and out-of-distribution regions.
The pervasive pixel distortions and background clutter in panoramic images often result in a dispersed pixel manifold. This dispersion adversely affects the performance of deep neural networks, which tend to map the unknown samples inside the known class clusters instead of an arbitrary distribution in the whole feature space. Consequently, this misalignment distorts the voting scores of the pixel manifold, particularly in out-of-distribution regions within known classes, ultimately leading to overconfident and unreliable predictions. CLIP, which is trained on large-scale vision-language data pairs, demonstrates robust cross-domain generalization capabilities. Moreover, its pretrained text encoder adeptly constructs the semantic feature distribution of known classes while maintaining consistency with the pixel space distribution. This accurate semantic feature distribution provides a reliable reference for pixel manifold learning. Leveraging this foundation, we propose a Bilevel Prompt Distribution Learning (BPDL) loss, which utilizes projected class and distribution prompts $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times C_{m}}$, to optimize the spatial distribution of per-pixel embeddings $F_{m}(x)$.